# DeepCausalMMM: A Deep Learning Framework for Marketing Mix Modeling with Causal Inference

### v1

@@
Marketing Mix Modeling (MMM) is a statistical technique used to estimate the impact of marketing activities on business outcomes such as sales, revenue, or customer visits. Traditional MMM approaches often rely on linear regression or Bayesian hierarchical models that assume independence between marketing channels and struggle to capture complex temporal dynamics and non-linear saturation effects (Hanssens et al. 2005; Ng et al. 2021).
@@
Key innovations include: (1) a data-driven design where hyperparameters and transformations (e.g., adstock decay, saturation curves) are learned or estimated from data with sensible defaults, rather than requiring fixed heuristics or manual specification, (2) multi-region modeling with both shared and region-specific parameters, (3) robust statistical methods including Huber loss and advanced regularization, (4) comprehensive response curve analysis for understanding channel saturation, and (5) an extensive visualization suite with 14+ interactive dashboards for business insights.
@@
Robyn (Meta) (Runge et al. 2024; contributors) 2024) uses Bayesian hyperparameter optimization with fixed adstock and saturation transformations (Adstock, Hill, Weibull). It provides budget optimization and is widely used in industry but requires manual specification of transformation types and does not model channel interdependencies.
LightweightMMM (Google) (contributors 2022) implements Bayesian MMM using JAX and Numpyro, offering probabilistic inference with flexible priors. It supports adstock effects and budget optimization but does not incorporate causal graph learning or deep learning for temporal dynamics.
@@
CausalMMM (Gong et al. 2024) introduces neural networks and graph learning to MMM, demonstrating the value of discovering channel interdependencies. However, it does not provide multi-region modeling, comprehensive response curve analysis, or the extensive visualization and analysis tools needed for practical deployment.
DeepCausalMMM advances the field by integrating: (1) GRU-based temporal modeling, (2) DAG-based structure learning (Zheng et al. 2018), (3) Hill equation response curves, (4) multi-region modeling, (5) robust statistical methods, (6) production-ready architecture, and (7) comprehensive visualization suite.
@@
  • Tests: 28 comprehensive tests with 100% pass rate
  • Visualizations: 14+ interactive Plotly dashboards exportable as HTML
@@
DeepCausalMMM ensures reproducible results through deterministic training with configurable random seeds, comprehensive test suite (28 tests), example notebooks, detailed documentation of hyperparameters, and version-controlled releases with semantic versioning.

### v2

@@
Marketing Mix Modeling (MMM) is a statistical technique used to estimate the impact of marketing activities on business outcomes such as sales, revenue, or customer visits. Traditional MMM approaches often rely on linear regression or Bayesian hierarchical models that assume independence between marketing channels and struggle to capture complex temporal dynamics and non-linear saturation effects (Hanssens et al. 2005; Chan & Perry 2017; Ng et al. 2021).
@@
Key features include: (1) a data-driven design where hyperparameters and transformations (e.g., adstock decay, saturation curves) are learned or estimated from data with sensible defaults, rather than requiring fixed heuristics or manual specification, (2) multi-region modeling with both shared and region-specific parameters, (3) robust statistical methods including Huber loss and advanced regularization, (4) comprehensive response curve analysis for understanding channel saturation.
@@
Robyn (Meta) (Runge et al. 2024; contributors) 2024) uses evolutionary hyperparameter search with fixed adstock and saturation transformations (Adstock, Hill, Weibull). It provides budget optimization and is widely used in industry but requires manual specification of transformation types and does not model channel interdependencies.
Meridian (Google) (Google Meridian Team 2025) is Google’s open-source Bayesian MMM framework featuring reach and frequency modeling, geo-level analysis, and experimental calibration. It employs causal inference with pre-specified causal graphs and the backdoor criterion.
@@
CausalMMM (Gong et al. 2024) introduces neural networks and graph learning to MMM, demonstrating the value of discovering channel interdependencies. However, it does not provide multi-region modeling, comprehensive response curve analysis.
DeepCausalMMM advances the field by integrating: (1) GRU-based temporal modeling, (2) DAG-based structure learning (Zheng et al. 2018), (3) Hill equation response curves, (4) multi-region modeling, (5) robust statistical methods, (6) production-ready architecture.
@@
  • Tests: Comprehensive unit and integration test suite in tests/ directory
@@
DeepCausalMMM ensures reproducible results through deterministic training with configurable random seeds, comprehensive test suite, example notebooks, detailed documentation of hyperparameters, and version-controlled releases with semantic versioning.